# A Survey on Multilingual Large Language Models: Corpora, Alignment, and Bias

### v2

@@ Abstract
Based on the foundation of Large Language Models (LLMs), Multilingual Large Language Models (MLLMs) have been developed to address the challenges of multilingual natural language processing tasks, hoping to achieve knowledge transfer from high-resource to low-resource languages. However, significant limitations and challenges still exist, such as language imbalance, multilingual alignment, and inherent bias. In this paper, we aim to provide a comprehensive analysis of MLLMs, delving deeply into discussions surrounding these critical issues. First of all, we start by presenting an overview of MLLMs, covering their evolution, key techniques, and multilingual capacities. Secondly, we explore widely utilized multilingual corpora for MLLMs’ training and multilingual datasets oriented for downstream tasks that are crucial for enhancing the cross-lingual capability of MLLMs. Thirdly, we survey the existing studies on multilingual representations and investigate whether the current MLLMs can learn a universal language representation. Fourthly, we discuss bias on MLLMs including its category and evaluation metrics, and summarize the existing debiasing techniques. Finally, we discuss existing challenges and point out promising research directions. By demonstrating these aspects, this paper aims to facilitate a deeper understanding of MLLMs and their potentiality in various domains.
Index Terms:
Large Language Model, Multilingual Large Language Model, Corpora, Alignment, Bias, Survey.I Introduction
@@
The emergence of Large Language Models (LLMs) has brought about a paradigm shift and revolution in the field of Natural Language Processing (NLP). This innovative approach trains a transformer-based model [1] on extensive volumes of data and then leverages fine-tuning or prompt learning to facilitate the model’s adaption to a wide variety of tasks. Based on the foundation of LLMs, large-scale Multilingual Large Language Models (MLLMs), such as mBERT [2], XLM [3], mT5 [4], BLOOM [5] and LLaMA [6], have been developed to tackle multilingual NLP tasks. MLLMs are pre-trained on a concatenation of texts in multiple languages with the hope that low-resource languages may benefit from high-resource languages due to linguistic similarities and shared representations inherent within language pairs.
@@
Compared to LLMs, MLLMs require larger multilingual corpora that cover more languages and diverse downstream tasks to ensure applicability and fairness across different languages. MLLMs are trained to understand and capture the structures and patterns of multiple languages. For instance, pre-trained on data from 104 languages, BLOOM supports 46 languages and 13 programming languages, covering the eight most widely spoken languages in the world [5]. Numerous MLLMs have been proposed in the past 5 years and the emergence of ChatGPT-like LLMs has accelerated the development of MLLMs. These MLLMs differ in the architecture (e.g., number of layers, parameters, etc), data used for pre-training (Wikipedia, Common Crawl, etc) and the number of languages involved (ranging from 12 to 110). However, it is uncertain how much cross-lingual transfer learning capability MLLMs have to support unseen languages or low-resource languages during pre-training. As a result, we first start by providing an overview of MLLMs in section II, which contains evolution, key techniques, and a detailed analysis of the multilingual capacities of MLLMs.
Despite the success of MLLMs, existing MLLMs still face numerous issues and challenges, which can be summarized as three aspects: corpora, alignment, and bias. As shown in figure 1, as input of MLLMs, corpora heavily influence the ability of MLLMs. On the one hand, unbalanced corpora and training lead to misalignment of MLLMs among different languages; On the other hand, inherent bias within corpora induces MLLMs to produce biased output. Therefore, our following discussion is presented around these three aspects.
@@
Firstly, MLLMs heavily rely on multilingual corpora to enhance their performance. For example, among the training corpus of ChatGPT, the English corpus accounts for 92.099%, and the Chinese only accounts for 0.16%, so its dialogues in the English context are much higher than those in other languages in terms of quality and speed. However, the size of available corpus resources for different languages varies greatly, and most of the existing labelled datasets are mainly focused on a few languages, limited effectiveness exists in zero-shot cross-lingual transfer from high-resource languages to languages that are unseen during training. Furthermore, MLLMs suffer from what Conneau et al. [7] call the curse of multilinguality: more languages lead to better cross-lingual performance on low-resource languages up until a point, after which the overall performance of MLLMs on monolingual and cross-lingual benchmarks will decrease. In conclusion, the scale, quality, and diversity of corpora have a significant impact on performance of MLLMs. Therefore, In section III, we present a survey of the multilingual corpora that representative MLLMs trained on, offering insights into their language distribution, data source and language coverage.
@@
Secondly, the success of MLLMs is their ability to achieve multilingual representation alignment from multiple languages. Aligning the representation of diverse languages acts as an integral part of NLP’s multilingual tasks and applications[8] and under-representation on low-resource languages leads to MLLMs’ poor performance on these languages. Inspired by the impressive performance of monolingual representation models like Word2vec [9] and GloVe [10], recent research has made great progress in multilingual representation. In section IV, we review previous research on multilingual word embeddings and classify them into three categories: static multilingual representation, contextual multilingual representation and combined multilingual representation. We also analyze the impact of various factors on multilingual alignment performance, including initial alignment solution, linearity of mapping function, language typological distance, and pre-training data and settings of MLLMs.
@@
Thirdly, MLLMs are prone to produce harmful outcomes and social bias [11] in part due to bias is naturally present in cross-cultural datasets and the design of MLLMs’ modeling processes [12]. Previous studies have explored bias in various NLP tasks and demographic groups. However, these studies are specific to English-based models [13, 14] and cannot be generalized to other languages. What are the types of bias in existing MLLMs? What are the main de-biasing techniques available for MLLMs? Does the removal of these biases affect the performance of the large language models? What are the existing bias evaluation datasets for MLLMs? These are very worthwhile research questions. This survey tries to answer these questions and offers valuable insights for bias on MLLMs.
Summarizing the above discussion, the contributions of this survey are as follows:
@@
We present an overview of MLLMs and analyze the the language imbalance challenge within MLLMs, their capacity to support low-resource languages and their potential for cross-lingual transfer learning.
@@
We provide an overview of the datasets and corpora utilized by existing MLLMs to offer a comprehensive insight into the language distribution within these training datasets.
@@
Our survey delves into bias within MLLMs, seeking to address essential questions such as identifying the types of bias present in current MLLMs, exploring prominent de-biasing techniques, summarizing available bias evaluation datasets for MLLMs.
II Overview of MLLMs
This section provides a brief overview of MLLMs, tracing their evolution from monolingual large language models (LLMs) to multilingual LLMs. It then illustrates the key techniques that contribute most to the success of MLLMs, as well as the multilingual capacities of MLLMs.
II-A Evolution of MLLMs
II-A1 Monolingual Evolution
The development of monolingual LLMs has made great progress in understanding and generating human languages. These models employ a transformer-based architecture to pre-train a large corpus of texts, followed by fine-tuning or prompt learning to enhance models’ performance on specific tasks or languages.
@@ II-A2 Multilingual Evolution
Following the success of monolingual BERT, multilingual BERT (mBERT) [2] was the first released MLLM by following the training procedure of BERT but on multilingual Wikipedia text corpora including 104 languages. Other MLLMs such as XLM-R [7], mBART[26] and mT5 [4] follow the step of mBERT, further explore the capacities and limitations of MLLMs across languages. Studies reveal that MLLMs work surprisingly well on cross-lingual tasks even without direct cross-lingual supervision like parallel or comparable data [27, 28].
@@
The further development trend has led to a tremendous increase in model parameters and data scale, resulting in enhancements that promote multilingual capabilities. For example, PaLM with 540 billion (540B) parameters yields impressive capabilities on the multilingual benchmarks by training on a mixture of multilingual versions of Wikipedia and dialogue data, including 124 languages [29]. With the early successful attempts, the autoregressive language modeling and prompt learning paradigm represented by GPT series have received much attention and follow-up from major companies and universities. Thus, more MLLMs (e.g., InstructGPT [21], LaMDA [30], OPT [31], BLOOM [5], LLaMA [6]) have been proposed to achieve breakthrough performance in a range of multi-step reasoning tasks over multiple languages. In addition to the GPT series and its derivative models, numerous other models have been proposed to boost the development of LLMs. Examples include GLM [32, 33], Vicuna [34], Gemini [35], and several others.
@@
The development of MLLMs has been guided by several tendencies: (1) parameter growth; (2) linguistic diversity; (3) multimodal unification. Regarding (1) MLLMs have been expanded to hundreds of billions of parameters or even trillions. Increasing the size of parameters brings clear benefits, such as alleviating hallucination phenomena heavily present in minor-parameter (e.g., 7B, 13B) models. However, there’s a limit to the amount of text data available online, and obtaining high-quality data is becoming increasingly challenging, which might slow down the parameter growth of MLLMs. Regarding (2), most high-resource languages belong to similar language families, thus sharing numerous linguistic features. Disregarding this diversity inevitably leads to poor generalizability and language-specific biases [36]. Recent work in MLLMs has focused on addressing this issue since low-resource and unseen languages still account for a large proportion of the world’s languages. Regarding (3), multimodal MLLMs are a growing focus of research, realizing a variety of specific real-world needs by unifying diverse types of modality (i.e., text, image, and speech). Additionally, current research aims to extend MLLMs to accommodate more modalities like web pages, heat maps, graphs, and tables, thereby increasing the model’s generality and applicability [37].
@@
Table I summarizes some representative MLLMs in recent years, divided into two categories: monolingual and multilingual, showing the evolution of MLLMs from multiple perspectives in chronological order of release.
@@
| PanGu-α [39] | Apr-21 | Huawei | 200B | 2K | 100GB | Decoder-only | - | LM | Open | Text |
@@
| Baichuan2 [50] | Sep-23 | BAICHUAN | 13B | 4K | 27.8GB | Decoder-only | Baichuan | LM | Open | Text |
@@
| Geimini-ultra [35] | Dec-23 | DeepMind | 300B | 32K | - | Decoder-only | - | LM | Closed | Text, Image |
| Gemini-pro [35] | Dec-23 | DeepMind | 100B | 32K | - | Decoder-only | - | LM | Closed | Text, Image |
@@
II-B Key Techniques of MLLMs
Transformer architecture, pre-training technique, and reinforcement learning with human feedback are the key techniques for MLLMs. In this section, we present the key idea of these techniques.
II-B1 Transformer Architecture
Transformer architecture, first introduced in 2017, has become the foundation of MLLMs owing to its suitability for parallel computing and flexibility for diverse model design. Transformer architecture consists of two main modules, an Encoder, and a Decoder, along with a self-attention mechanism within the modules. The encoder, using stacked multi-head self-attention layers, encodes the input sequence and generates latent representations. In contrast, the decoder employs cross-attention to utilize the encoder’s latent representations, attending to them while autoregressively generating the target sequence [53].
@@ II-B1 Transformer Architecture
Encoder-only (e.g., GLM ): MLLMs with encoder-only architecture can effectively handle long-range dependencies within the input sequences, making them well-suited for the analysis and classification of textual content, including tasks like sentiment analysis and named entity recognition.
@@
Decoder-only (e.g., PaLM): MLLMs with decoder-only architecture are mainly designed to generate sequences of language texts. They predict the next token based on contextual information from the current and preceding steps.
@@
Encoder-decoder hybrid (e.g., mT5): MLLMs with encoder-decoder architecture enable themself to process sequential data and generate accurate and coherent outputs that excel in tasks such as text generation, and summarization.
II-B2 Pre-training Technique
@@
Pre-training technique aims to learn universal language representations from billion-scale unlabeled corpus (e.g., Wikipedia, Webpages, News, etc.) and then initializes the parameters of the Transformer-based MLLMs. This approach reduces the reliance on massive parallel corpus, helping MLLMs generate similar representations in a common vector space for similar sentences and words (or words in similar context) across languages[54].
@@
The benefits of pre-training technique can be attributed to two key factors: Paradigm and Task. Pre-training paradigms have been proposed to capture linguistic patterns in the training data and adapt MLLMs to downstream tasks, including “pre-training + fine-tuning” and “pre-training + prompting”. The former representative models are BERT [2], GPT-2 [19], while the latter presentative models like GPT-3 [20]. Pre-training tasks improve the ability of MLLMs to encode and generate coherent multilingual text.
@@
In learning the universal representation of language, pre-training tasks play a crucial role and the widely used pre-training tasks include probabilistic language modeling (LM), masked language modeling (MLM), next sentence prediction (NSP), and Denoising autoencoder (DAE). Probabilistic LM is a fundamental task in NLP, estimating the probability distribution of sequences of words in a language. In practice, LM typically involves auto-regressive LM or unidirectional LM. MLM has emerged as a novel pre-training task to overcome the drawback of the standard unidirectional LM. By masking certain tokens in a sequence and predicting them based on context, MLM encourages models to learn bidirectional representations, capturing dependencies from both left and right contexts. Punctuations are the natural separators of text data. So, it is reasonable to construct pre-training methods by utilizing them. NSP is just a great example of this. NSP encourages the model to understand the contextual coherence and relationships between sentences. DAE takes a partially corrupted input and aims to recover the original undistorted input. Specific to language, a sequence-to-sequence model, such as the standard Transformer, is used to reconstruct the original text. Eq.1 summarize loss function of these pre-training tasks, where denotes a sequence of tokens [55].
II-B3 Reinforcement Learning with Human Feedback

### v3

@@ Abstract
Based on the foundation of Large Language Models (LLMs), Multilingual LLMs (MLLMs) have been developed to address the challenges faced in multilingual natural language processing, hoping to achieve knowledge transfer from high-resource languages to low-resource languages. However, significant limitations and challenges still exist, such as language imbalance, multilingual alignment, and inherent bias. In this paper, we aim to provide a comprehensive analysis of MLLMs, delving deeply into discussions surrounding these critical issues. First of all, we start by presenting an overview of MLLMs, covering their evolutions, key techniques, and multilingual capacities. Secondly, we explore the multilingual training corpora of MLLMs and the multilingual datasets oriented for downstream tasks that are crucial to enhance the cross-lingual capability of MLLMs. Thirdly, we survey the state-of-the-art studies of multilingual representations and investigate whether the current MLLMs can learn a universal language representation. Fourthly, we discuss bias on MLLMs, including its categories, evaluation metrics, and debiasing techniques. Finally, we discuss existing challenges and point out promising research directions of MLLMs.
keywords:
[…]
UTF8gbsn
1 Introduction
@@
The rapid development of Large Language Models (LLMs) has brought about a paradigm shift and revolution in the field of Natural Language Processing (NLP). This innovative approach trains a transformer-based model [1] on extensive volumes of data and then leverages fine-tuning or prompt learning to facilitate the model’s adaption to a wide variety of tasks. Based on the foundation of LLMs, large-scale Multilingual MLLMs (MLLMs), such as mBERT [2], XLM [3], mT5 [4], BLOOM [5] and LLaMA [6], have been developed to tackle multilingual NLP tasks. MLLMs are pre-trained on a concatenation of texts in multiple languages with the hope that low-resource languages may benefit from high-resource languages due to linguistic similarities and shared representations inherent within language pairs.
@@
Compared to LLMs, MLLMs require larger multilingual corpora that cover more languages to ensure applicability and fairness across different languages in downstream tasks. MLLMs are trained to understand and capture the structures and patterns of multiple languages. For instance, pre-trained on data from 104 languages, BLOOM supports 46 languages, covering the eight most widely spoken languages in the world [5]. Numerous MLLMs have been proposed in the past 5 years, which differ in the architecture (e.g., number of layers, parameters, etc), data used for pre-training (Wikipedia, Common Crawl, etc), and the number of languages involved (ranging from 12 to 110). However, it is uncertain how much cross-lingual transfer learning capability MLLMs have to support unseen languages or low-resource languages during pre-training. As a result, section 2 first starts by providing an overview of MLLMs, which contains key evolutions, techniques, and a detailed analysis of MLLMs’ multilingual capacities. Despite the success of MLLMs, existing MLLMs still face numerous issues and challenges, which can be summarized as three aspects: corpora, alignment, and bias. As shown in Figure 1, training corpora of MLLMs heavily influence their capability. On the one hand, the unbalanced corpora and training lead to misalignment of MLLMs among different languages; On the other hand, inherent bias within corpora induces MLLMs to produce biased output. Therefore, this paper focuses discussion around the three aspects of corpora, alignment, and bias.
@@
Firstly, MLLMs heavily rely on multilingual corpora to enhance their performance. For example, among the training corpus of ChatGPT, the English corpus accounts for 92.099%, and the Chinese only accounts for 0.16%, so its dialogues in the English context are much higher than those in other languages in terms of quality and speed. However, the size of available corpus resources for different languages varies greatly, and most of the existing annotated datasets are mainly focused on a few languages, limited the cross-lingual transfer effectiveness from high-resource languages to languages that are unseen during training. Furthermore, MLLMs suffer from what Conneau et al. [7] call the curse of multilinguality: more languages lead to better cross-lingual performance on low-resource languages up until a point, after which the overall performance of MLLMs on monolingual and cross-lingual benchmarks will decrease. In conclusion, the scale, quality, and diversity of corpora have a significant impact on the performance of MLLMs. Therefore, section 3 presents a survey of the representative multilingual training corpora of MLLMs, offering insights into their language distribution, data source, and language coverage.
@@
Secondly, MLLMs still struggle to learn a universal language representation for diverse languages. The language misalignment issues exist. Aligning the representation of diverse languages acts as an integral part of NLP’s multilingual tasks and applications[8], and under-representation of low-resource languages leads to MLLMs’ poor performance on these languages. Inspired by the impressive performance of monolingual representation models like Word2vec [9] and GloVe [10], recent research has made great progress in multilingual representation. In section 4, we review previous research on multilingual representations and classify them into three categories: static multilingual representation, contextual multilingual representation, and combined multilingual representation. We also analyze the impact of various factors on multilingual alignment performance, including initial alignment solution, linearity of mapping function, language typological distance, and pre-training data and settings of MLLMs.
@@
Thirdly, MLLMs are prone to produce harmful outcomes and social bias [11] in part due to bias is naturally present in cross-cultural datasets and the design of MLLMs’ modeling processes [12]. Previous studies have explored bias in various NLP tasks and demographic groups,but are largely specific to English-based models [13, 14], which cannot be generalized to other languages. What are the types of bias in existing MLLMs? What are the main de-biasing techniques available for MLLMs? Does the removal of these biases affect LLMs’ performance? What are the existing bias evaluation datasets for MLLMs? These are very worthwhile research questions. This survey tries to answer these questions and offers valuable insights for bias on MLLMs.
The contributions of this survey are as follows:
@@
We present an overview of MLLMs and analyze the language imbalance challenge within MLLMs, their capacity to support low-resource languages and their potential for cross-lingual transfer learning.
@@
We provide an overview of the multilingual datasets and corpora utilized by existing MLLMs, offering a comprehensive insight into the language distribution within these training corpora.
@@
Our survey delves into bias within MLLMs, seeking to address essential questions such as identifying the types of bias present in current MLLMs, exploring prominent de-biasing techniques, and summarizing available bias evaluation datasets for MLLMs.
2 Overview of MLLMs
This section provides a brief overview of MLLMs, tracing their evolution from monolingual LLMs to multilingual LLMs, as dispicted in Figure 2. It then illustrates the key techniques that contribute most to the success of MLLMs, as well as the multilingual capacities of MLLMs.
2.1 Evolution of MLLMs
2.1.1 Monolingual Evolution
The development of monolingual LLMs has made great progress in understanding and generating human languages. These models employ a transformer-based architecture that is first pre-trained on a large corpus of texts, followed by fine-tuning or prompt learning to enhance models’ performance on specific tasks or languages.
@@ 2.1.2 Multilingual Evolution
Following the success of monolingual BERT, multilingual BERT (mBERT) [2] was the first released MLLM by following the training procedure of BERT but on multilingual Wikipedia text corpora including 104 languages. Other MLLMs such as XLM-R [7], mBART[26] and mT5 [4] follow the step of mBERT, further exploring the capacities and limitations of MLLMs across languages. Studies reveal that MLLMs work surprisingly well on cross-lingual tasks even without direct cross-lingual supervision like parallel or comparable data [27, 28].
@@
The further development trend has led to a tremendous increase in model parameters and data scale, resulting in enhancements that promote multilingual capabilities. For example, PaLM with 540 billion (540B) parameters yields impressive capabilities on the multilingual benchmarks by training on a mixture of multilingual versions of Wikipedia and dialogue data, including 124 languages [29]. With the early successful attempts, the autoregressive language modeling and prompt learning paradigm represented by the GPT series has received much attention and follow-up from major companies and universities. Thus, more MLLMs (e.g., InstructGPT [21], LaMDA [30], OPT [31], BLOOM [5], LLaMA [6]) have been proposed to achieve breakthrough performance in a range of multi-step reasoning tasks over multiple languages. In addition to the GPT series and its derivative models, numerous other models have been proposed to boost the development of LLMs. Examples include GLM [32, 33], Vicuna [34], Gemini [35], and several others.
@@
The development of MLLMs has been guided by several tendencies: (1) parameter growth; (2) linguistic diversity; (3) multimodal unification. Regarding (1), MLLMs have expanded to hundreds of billions of parameters or even trillions. Increasing the size of parameters brings clear benefits, such as alleviating hallucination phenomena heavily present in minor-parameter (e.g., 7B, 13B) models. However, there’s a limit to the amount of text data available online, and obtaining high-quality data is becoming increasingly challenging, which might slow down the parameter growth of MLLMs. Regarding (2), most high-resource languages belong to similar language families, thus sharing numerous linguistic features. Disregarding this diversity inevitably leads to poor generalizability and language-specific biases [36]. Recent work in MLLMs has focused on addressing this issue as low-resource and unseen languages still account for a large proportion of the world’s languages. Regarding (3), multimodal MLLMs are a growing focus of research, realizing a variety of specific real-world needs by unifying diverse types of modalities (i.e., text, image, and speech). Additionally, current research aims to extend MLLMs to accommodate more modalities like web pages, heat maps, graphs, and tables, thereby increasing the model’s generality and applicability [37].
@@
Table 1 summarizes the representative MLLMs in recent years, divided into two categories: monolingual and multilingual, showing the evolution of MLLMs from multiple perspectives in chronological order of release.
@@
| Gemini-ultra [35] | Dec-23 | DeepMind | 300B | 32K | - | Decoder-only | - | LM | Closed | Text, Image |
@@
| Claude 3 [51] | Mar-24 | Anthropic | - | 200K | - | Decoder-only | Claude | LM | Closed | Text, Image |
| LLaMA 3 [52] | Apr-24 | Meta | 70B | 8K | 140GB | Decoder-only | LLaMA | LM | Open | Text |
2.2 Key Techniques of MLLMs
Transformer architecture, pre-training technique, and reinforcement learning with human feedback are the key techniques for MLLMs. In this section, we will present the key ideas behind these techniques.
2.2.1 Transformer Architecture
Transformer architecture, first introduced in 2017, has become the foundation of MLLMs owing to its suitability for parallel computing and flexibility for diverse model design. Transformer architecture consists of two main modules, Encoder and Decoder, along with a self-attention mechanism within each module. The encoder, using stacked multi-head self-attention layers, encodes the input sequence and generates latent representations. In contrast, the decoder employs cross-attention to utilize the encoder’s latent representations, attending to them while autoregressively generating the target sequence [53].
@@ 2.2.1 Transformer Architecture
Encoder-only (e.g., BERT ): MLLMs with encoder-only architecture can effectively handle long-range dependencies within the input sequences, making them well-suited for the analysis and classification of textual content, including tasks like sentiment analysis and named entity recognition.
@@
Decoder-only (e.g., GPT): MLLMs with decoder-only architecture are mainly designed to generate sequences of language texts. They predict the next token based on contextual information from the current and preceding steps.
@@
Encoder-decoder hybrid (e.g., GLM): MLLMs with encoder-decoder architecture enable themselves to process sequential data and generate accurate and coherent outputs that excel in tasks such as text generation, and summarization.
2.2.2 Pre-training Technique
@@
Pre-training technique aims to learn universal language representations from billion-scale unlabeled corpora (e.g., Wikipedia, webpages, news, etc.) and then initializes the parameters of the Transformer-based MLLMs. This approach reduces the reliance on massive parallel corpora, helping MLLMs generate similar representations in a common vector space for similar sentences and words (or words in similar context) across languages[54].
@@
The benefits of pre-training technique can be attributed to two key factors: Paradigm and Task. Pre-training paradigms have been proposed to capture linguistic patterns in the training data and adapt MLLMs to downstream tasks, including “pre-training + fine-tuning" and “pre-training + prompting". The former representative models are BERT [2], GPT-2 [19], while the latter presentative models like GPT-3 [20]. Pre-training tasks improve the ability of MLLMs to encode and generate coherent multilingual text.
@@
When learning the universal representation of language, pre-training tasks play a crucial role and the widely used pre-training tasks include probabilistic language modeling (LM), masked language modeling (MLM), next sentence prediction (NSP), and Denoising autoencoder (DAE). Probabilistic LM is a fundamental task in NLP, estimating the probability distribution of sequences of words in a language. In practice, LM typically involves auto-regressive LM or unidirectional LM. MLM has emerged as a novel pre-training task to overcome the drawback of the standard unidirectional LM. By masking certain tokens in a sequence and predicting them based on context, MLM encourages models to learn bidirectional representations, capturing dependencies from both left and right contexts. Punctuations are the natural separators of text data. So, it is reasonable to construct pre-training methods by utilizing them. NSP is just a great example of this. NSP encourages the model to understand the contextual coherence and relationships between sentences. DAE takes a partially corrupted input and aims to recover the original undistorted input. Specific to language, a sequence-to-sequence model, such as the standard Transformer, is used to reconstruct the original text. Eq.1 summarize loss function of these pre-training tasks, where denotes a sequence of tokens [55].
where the symbols and their definitions are listed in Table 2.
| Symbol | Description |
| x | |
| m(x) | The masked tokens |
| The rest tokens | |
| Random perturbation text | |
| y | Target label |
| If x and y are continuous segments from corpus, | |
| The prompt | |
| The better model response | |
| The worse model response | |
| The output of reward model | |
| Weight in policy optimization network | |
| The objective function | |
| The gradient of objective function |
2.2.3 Reinforcement Learning with Human Feedback